OpenStudy (anonymous):

What is the geometric mean of 3 and 21?

OpenStudy (anonymous):

3/x = x/21 or x^2 = 3(21) or x^2 = 63 or x = sqrt(63) Do we need to simplify the radical?

OpenStudy (anonymous):

sqrt ( 3*21)

OpenStudy (anonymous):

Its a multiple choice question It gives 3 radical 7 ; 7 radical 3; 31.5 ;63

OpenStudy (anonymous):

x = sqrt (9 * 7) or \[x = 3\sqrt{7}\]

OpenStudy (anonymous):

Do you understand or just need the answer?

OpenStudy (anonymous):

no i wasnt sure how to get the answer

OpenStudy (anonymous):

okay so we need to simplify radicals. This might come in handy later ...

OpenStudy (anonymous):

okay